Consider an electron in the presence of an externally applied magnetic field. If the magnetic force on the electron points to the left and the magnetic field points out of the page/screen, what is the direction of the electron's velocity?
Answer #1

Since F = q(V × B) = - e (V×B)

That means F , V and B are perpendicular to each other.

Here F = - Fi^ (left)

B = Bk^ ( out of page)

From right hand rule , the direction of electron is opposite to the direction of positive charge .

Hence direction of velocity will be along negative y axis that means downward .

V = - Vj^